How to find the maximum Pythagorean triplets if altitude is 40
1
Expert's answer
2019-12-27T08:43:03-0500

Formulas for obtaining Pythagorean triples:

a = 2m +1 , b = 2m(m+1) = 2m2 + 2m , c = 2m2 + 2m + 1.

If m = 4,

a = 2*4 +1= 9,

b = = 2*42 + 2*4 = 40 ,

c = 2*42 + 2*4 + 1 = 41.

92 + 402 = 412

9, 40, 41